A certain piggy bank contains only pennies and quarters. It contains four times as many pennies as quarters . If the bank contains 440 coins, how many are quarters?
Answer by ikleyn(52807)

.

x quarters and 4x pennies, in all


    x + 4x = 440   coins.


    5x     = 440

     x     = 440/5 = 88.


ANSWER.  88 quarters and 4*88 = 352 pennies.